OpenStudy (nuccioreggie):

58=12+2× Subtract 12 from both sides

OpenStudy (anonymous):

\[58=12~+~2x\]\[58 - 12 = 12 - 12 + 2x\]\[46 = 2x\]\[\frac{46}{2} = \frac{2x}{2}\]

OpenStudy (nuccioreggie):

@kelliegirl33

OpenStudy (anonymous):

Can you finish it?

OpenStudy (nuccioreggie):

Is the answer for my question 0 ?

OpenStudy (anonymous):

no... 46/2 = 2x/2 -- the 2's on the right side cancel......so just divide 46 by 2 and you will have found x

OpenStudy (anonymous):

look at it this way.. 46/2 = (2/2)x 46/2 = 1x or just 46/2 = x......now just divide 46/2

OpenStudy (nuccioreggie):

So my answer for this question is 23x ??

OpenStudy (anonymous):

your answer would be : x = 23

OpenStudy (nuccioreggie):

Ooooook thnk you

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

\[58=12~+~2x\]\[58 - 12 = 12 - 12 + 2x\]\[46 = 2x\]\[\frac{46}{2} = \frac{2x}{2}\]\[23 = x\]

OpenStudy (nuccioreggie):

So now can you sdivideboth sides by 2

OpenStudy (anonymous):

You divide both sides by two to get `x` alone on one side.

OpenStudy (anonymous):

exactly ^^

OpenStudy (nuccioreggie):

Ok so how whats the answer going to be?

OpenStudy (anonymous):

The answer is \(23 = x\).

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @nuccioreggie So my answer for this question is 23x ?? \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @kelliegirl33 your answer would be : x = 23 \(\color{blue}{\text{End of Quote}}\) You and KellieGirl have both already said it. :)

OpenStudy (nuccioreggie):

So its 23=x if i divide 2 on both sides olso right?

OpenStudy (nuccioreggie):

??

OpenStudy (nuccioreggie):

@kelliegirl33

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (anonymous):

Correct.